Solving for x and y, what are the variables in the following equations solving systems of equations?
4x+y=32
−3x−y=−26

Answer by macston(5194)

4x+y=32
-3x-y=-26 Add equations.
x=6 Use x in one of the equations to solve for y.
4x+y=32
4(6)+y=32
24+y=32
y=8 ANSWER: x=6 and y=8
CHECK:
Use other equation to check.
-3x-y=-26
-3(6)-8=-26
-18-8=-26
-26=-26